# STRelay: A Universal Spatio-Temporal Relaying Framework for Location Prediction with Future Spatiotemporal Contexts

### v1

@@ Introduction
Multi-Task Location Prediction. After the overall context embedding is generated, it is then fed into a multi-layer perceptron (MLP) followed by a softmax function to predict the next location:
@@
Experiments
Table 1: Datasets Statistics
@@
Table 2: Overall location prediction performance
Istanbul Tokyo Singapore Moscow
Methods Acc@10 NDCG@10 MRR Acc@10 NDCG@10 MRR Acc@10 NDCG@10 MRR Acc@10 NDCG@10 MRR
BPR 0.3816 0.2500 0.1528 0.3215 0.2121 0.1317 0.3193 0.2069 0.1176 0.2093 0.1273 0.0820
Caser 0.3751 0.2649 0.2396 0.5000 0.3275 0.2830 0.3091 0.1953 0.1600 0.3974 0.2579 0.2301
STGCN 0.3949 0.2721 0.2428 0.5306 0.3422 0.2927 0.3691 0.2353 0.2138 0.4237 0.2810 0.2437
DeepMove 0.3368 0.2141 0.1859 0.5107 0.3338 0.2884 0.3871 0.2360 0.2009 0.3790 0.2322 0.1967
STAN 0.3716 0.2467 0.2171 0.4890 0.3147 0.2702 0.3855 0.2382 0.2040 0.3993 0.2474 0.2102
GETNext 0.3816 0.2536 0.2230 0.5197 0.3408 0.2948 0.4008 0.2564 0.2217 0.4302 0.2831 0.2456
STGN 0.3936 0.2704 0.2410 0.5342 0.3529 0.3056 0.3538 0.2209 0.1905 0.4157 0.2755 0.2392
Flashback 0.4139 0.2847 0.2557 0.5459 0.3640 0.3119 0.3815 0.2425 0.2092 0.4410 0.3007 0.2633
SNPM 0.4125 0.2849 0.2527 0.5470 0.3607 0.3109 0.3993 0.2644 0.2164 0.4562 0.3118 0.2738
Graph-Flashback 0.4173 0.2925 0.2578 0.5548 0.3647 0.3154 0.4068 0.2648 0.2231 0.4645 0.3197 0.2765
STRelay+STGN 0.4316 0.2921 0.2576 0.5569 0.3709 0.3232 0.3748 0.2350 0.2024 0.4633 0.3076 0.2669
STRelay+Flashback 0.4473 0.3072 0.2721 0.5743 0.3859 0.3377 0.3978 0.2523 0.2175 0.4740 0.3177 0.2756
STRelay+SNPM 0.4484 0.3070 0.2723 0.5786 0.3837 0.3329 0.4172 0.2663 0.2259 0.4944 0.3258 0.2877
STRelay+Graph-Flashback 0.4525 0.3101 0.2743 0.5835 0.3862 0.3381 0.4251 0.2728 0.2338 0.4970 0.3360 0.2915
Experiment Setup
@@
Dataset. We evaluate our STRelay using four widely used real-world datasets: Istanbul, Tokyo, Singapore and Moscow. They are collected from April 2012 to February 2013 (Yang et al. 2019). Each check-in consists of the user ID, POI ID, latitude, longitude, and timestamp. We follow the data pre-processing steps in (Yang et al. 2020). The initial 80% of the data (in chronological order) is used for training, and the remaining data for testing. Table 1 summarizes the details of the four processed datasets.
Baselines. We choose the following models for next location prediction as baselines: BPR (Rendle et al. 2012), Caser (Tang and Wang 2018), STGN (Zhao et al. 2019), STGCN (Zhao et al. 2019), DeepMove (Feng et al. 2018), STAN (Luo, Liu, and Liu 2021), GETNext (Yang, Liu, and Zhao 2022), Flashback (Yang et al. 2020), SNPM (Yin et al. 2023), and Graph-Flashback (Rao et al. 2022). Please refer to Appendix for the details of baselines.
@@
Among these baselines, we instantiate our STRelay with four techniques: STGN, Flashback, SNPM, and Graph-Flashback to validate its effectiveness.
@@
Evaluation Metrics. We adopt three evaluation metrics commonly used in prior work: Mean Reciprocal Rank (MRR), average Accuracy@K (Acc@K), and Normalized Discounted Cumulative Gain@K (NDCG@K) to evaluate the performance. In our experiments, we adopt the commonly used values $K=\{5,10\}$. Please refer to the Appendix for the details of the metrics.
Overall Performance Comparison
Table 2 shows the overall performance on all four datasets (due to limited space, we report the results of Acc@5 and NDCG@5 in Section Additional Results on Overall Performance Comparison in the Appendix), and we highlight the best-performing result on each task and for each dataset. We observe that STRelay+Graph-Flashback consistently outperforms all baselines, achieving 4.89%-8.47% improvement on average over the best-performing baselines across four datasets. Moreover, DeepMove, STAN and GETNext consider future temporal information but still yield suboptimal performance compared to STRelay enhanced baselines.
Notably, we observe that our STRelay framework consistently improves the location prediction performance of the corresponding base prediction methods STGN, Flashback, SNPM, and Graph-Flashback on all datasets, which demonstrates the universal effectiveness of STRelay. STRelay boosts the performance of the above four baselines with an average improvement of 3.19%-11.56% across different datasets in predicting the next location.
Ablation Study
We consider three variants to quantify their impact on location prediction performance and demonstrate their utility. 1) w/o spatial is a variant that only considers future temporal contexts for location prediction. 2) w/o temporal is a variant that only considers future spatial contexts for location prediction. 3) w/o relaying is a variant that models future temporal and spatial contexts not in a relaying, but in a parallel manner.
The results are shown in Table 3. First, the complete STRealy framework significantly outperforms both the spatial-ablated variant (w/o spatial) and temporal-ablated variant (w/o temporal) by 2.89% and 2.51% on average across four base models and four datasets, respectively, highlighting the essential contributions of modeling both temporal and spatial context. Moreover, we notice that the complete STRelay framework outperforms the w/o relaying variant by 1.49%-2.92% performance improvement across four base models and datasets, which demonstrates that our learning in a relaying manner indeed boosts performance by flexibly capturing the inherent relationship between temporal and spatial contexts. Note that due to limited space, we report the results of Acc@5 and NDCG@5 in Section Additional Results on Ablation Study in the Appendix.
(a) Location types
(b) User groups
Figure 4: Performance improvement on different location types and user groups
Table 3: Ablation Study
Istanbul Tokyo Singapore Moscow
Methods Acc@10 NDCG@10 MRR Acc@10 NDCG@10 MRR Acc@10 NDCG@10 MRR Acc@10 NDCG@10 MRR
STRelay+STGN 0.4316 0.2921 0.2576 0.5569 0.3709 0.3232 0.3748 0.2347 0.2024 0.4633 0.3076 0.2669
w/o spatial 0.4050 0.2790 0.2491 0.5373 0.3599 0.3138 0.3597 0.2286 0.1980 0.4361 0.2901 0.2517
w/o temporal 0.4013 0.2763 0.2468 0.5381 0.3609 0.3148 0.3550 0.2258 0.1958 0.4443 0.2966 0.2574
w/o relaying 0.4095 0.2819 0.2515 0.5437 0.3633 0.3161 0.3659 0.2330 0.2017 0.4420 0.2948 0.2562
STRelay+Flashback 0.4473 0.3073 0.2721 0.5743 0.3859 0.3377 0.3978 0.2523 0.2175 0.4861 0.3251 0.2816
w/o spatial 0.4340 0.2999 0.2668 0.5559 0.3672 0.3171 0.3902 0.2491 0.2155 0.4688 0.3137 0.2723
w/o temporal 0.4366 0.3012 0.2677 0.5586 0.3695 0.3194 0.3939 0.2508 0.2167 0.4764 0.3212 0.2794
w/o relaying 0.4404 0.3029 0.2688 0.5651 0.3807 0.3316 0.3942 0.2511 0.2173 0.4707 0.3156 0.2740
STRelay+SNPM 0.4484 0.3070 0.2723 0.5786 0.3837 0.3329 0.4172 0.2663 0.2259 0.4944 0.3258 0.2877
w/o spatial 0.4364 0.3011 0.2677 0.5695 0.3781 0.3271 0.4040 0.2543 0.2185 0.4827 0.3270 0.2851
w/o temporal 0.4402 0.3032 0.2691 0.5715 0.3747 0.3222 0.4115 0.2575 0.2206 0.4877 0.3284 0.2853
w/o relaying 0.4403 0.3018 0.2674 0.5806 0.3851 0.3327 0.4015 0.2536 0.2182 0.4899 0.3286 0.2859
STRelay+Graph-Flashback 0.4525 0.3101 0.2743 0.5835 0.3862 0.3381 0.4251 0.2728 0.2338 0.4993 0.3360 0.2915
w/o spatial 0.4432 0.3052 0.2709 0.5763 0.3844 0.3333 0.4107 0.2594 0.2231 0.4934 0.3331 0.2898
w/o temporal 0.4453 0.3068 0.2723 0.5773 0.3805 0.3279 0.4110 0.2586 0.2220 0.4986 0.3349 0.2903
w/o relaying 0.4368 0.3009 0.2672 0.5741 0.3819 0.3305 0.4216 0.2685 0.2312 0.4827 0.3256 0.2832
Performance Improvement Analysis on Location Types and User Groups
First, to study the performance improvement across different location types, we classify locations into corresponding activity categories following previous work (Yang et al. 2019). Specifically, we adopt a binary classification scheme consisting of: leisure-related locations (e.g., Outdoors and Entertainment) and daily-routine-related locations (e.g., College and Residence). Figure 4(a) shows the results of the percentage of performance improvement of Acc@10 in predicting these two types of locations on the four datasets. We observe that the performance improvements stemming from leisure-related locations are consistently and significantly higher than those from daily-routine-related locations by 2.68%-4.25%, which demonstrates that modeling future spatiotemporal contexts is particularly beneficial for entertainment-related activities, which are more uncertain than daily-routine-related activities.
Second, we categorize users into two distinct groups based on their radius of gyration (Feng et al. 2020; Deng et al. 2025a), which is calculated as the root mean squared distance of all locations in a trajectory from the geographical center of the trajectory, and obtain two groups: long-distance preferred users (top 50%) and short-distance preferred users (bottom 50%). We report the performance improvements on these two user groups. As shown in Figure 4(b), both user groups exhibit performance improvements in Acc@10 across all four datasets, with long-distance preferred users consistently demonstrating larger improvement (0.45%-1.06% higher than short-distance preferred users). These results indicate that modeling future spatiotemporal contexts is more effective in capturing the mobility patterns of users who prefer traveling longer distances, which often implies higher uncertainty for next locations.
Figure 5: Impact of the granularity of temporal and spatial intervals on the Istanbul dataset
Impact of Spatiotemporal Context Granularity
In this section, we use STRelay with Graph-Flashback to investigate the influence of the granularity of future temporal and spatial intervals on the performance (we report results of Acc@10 here). Results are shown in Figure 5. Specifically, to investigate the impact of temporal interval granularity, we train the model with future time intervals ranging from 0.5 to 24 hours. We observe the performance decreases as temporal granularity becomes coarser, with 1h granularity achieving the best results. For spatial interval granularity, we use spatial intervals from 0.5 km to 30 km. The results show a similar trend, where performance drops as spatial granularity coarsens, with 1km granularity achieving the best results. Thus, we choose 1h and 1km as temporal and spatial granularities. (Similar trends are observed in the other three datasets; please refer to the Appendix for the results.)
Conclusion
In this paper, we propose STRelay, a universal framework designed to explicitly model the future spatiotemporal contexts in a relaying manner; it can be flexibly integrated with different location prediction models to boost the prediction performance. Experimental results show that STRelay consistently boosts the performance of state-of-the-art baselines with an average improvement of 3.19%-11.56% across four datasets, demonstrating the universal effectiveness of STRelay. Moreover, we find that the future spatiotemporal contexts are particularly helpful for entertainment-related locations and also for user groups who prefer traveling longer distances. The performance gain on such non-daily-routine activities, which often suffer from higher uncertainty, is indeed complementary to the base location prediction models that excel at modeling regular daily routine patterns.
Appendix
Table 4: Additional results of location prediction performance
Istanbul Tokyo Singapore Moscow
Methods Acc@5 NDCG@5 Acc@5 NDCG@5 Acc@5 NDCG@5 Acc@5 NDCG@5
BPR 0.2532 0.1821 0.2133 0.1565 0.2077 0.1466 0.1362 0.0937
Caser 0.3179 0.2464 0.4115 0.2988 0.2392 0.1726 0.3239 0.2341
STGCN 0.3312 0.2514 0.4331 0.3104 0.2917 0.2102 0.3552 0.2587
DeepMove 0.2633 0.1903 0.4127 0.3020 0.2923 0.2053 0.2905 0.2035
STAN 0.2633 0.1903 0.4127 0.3020 0.2923 0.2053 0.2905 0.2035
GETNext 0.3107 0.2306 0.4214 0.3089 0.3171 0.2293 0.3528 0.2580
STGN 0.3292 0.2495 0.4415 0.3228 0.2736 0.1950 0.3475 0.2533
Flashback 0.3475 0.2632 0.4533 0.3339 0.2995 0.2159 0.3780 0.2802
SNPM 0.3458 0.2633 0.4538 0.3305 0.3138 0.2368 0.3913 0.2906
Graph-Flashback 0.3521 0.2711 0.4589 0.3346 0.3218 0.2376 0.3986 0.2967
STRelay+STGN 0.3604 0.2690 0.4625 0.3412 0.2914 0.2080 0.3841 0.2819
STRelay+Flashback 0.3779 0.2847 0.4803 0.3552 0.3116 0.2243 0.4009 0.2939
STRelay+SNPM 0.3781 0.2846 0.4801 0.3521 0.3279 0.2382 0.4167 0.3017
STRelay+Graph-Flashback 0.3807 0.2873 0.4863 0.3555 0.3348 0.2433 0.4217 0.3113
Baselines
We choose the following models for location prediction as baselines: BPR (Rendle et al. 2012) learns user preferences on POIs by minimizing a pairwise ranking loss. Caser (Tang and Wang 2018) embeds POIs in user interaction history as images, using convolutional filters to capture both local and global sequential dependencies. STGN (Zhao et al. 2019): add additional gates controlled by the spatiotemporal distances between successive check-ins to LSTM, while STGCN (Zhao et al. 2019)is a variant of STGN with coupled input and forget gates for improved efficiency. Note that we add user preference modeling in our experiments for these two methods. DeepMove (Feng et al. 2018) adds an attention mechanism to GRU for location prediction over sparse mobility traces. STAN (Luo, Liu, and Liu 2021) Models relative spatial-temporal information among POIs with a bi-layer attention architecture. GETNext (Yang, Liu, and Zhao 2022) utilizes a global trajectory flow graph to enhance the prediction of the next POI, alongside proposing a GCN model for generating effective POI embeddings. Flashback (Yang et al. 2020) is an RNN-based framework that explicitly leverages high-order spatiotemporal distance information to identify informative past hidden states. SNPM (Yin et al. 2023) represents POI relationships through knowledge graph embeddings and aggregates contextual information from neighboring nodes. Graph-Flashback (Rao et al. 2022) constructs a spatial-temporal knowledge graph for POI representation learning.
The Details of Evaluation Metrics
In this section, we elaborate on the details of evaluation metrics. The definitions of Mean Reciprocal Rank (MRR), average Accuracy@K (Acc@K), and Normalized Discounted Cumulative Gain@K (NDCG@K) are as follows:
@@ Conclusion
Table 5: Additional results of ablation study
Istanbul Tokyo Singapore Moscow
Methods Acc@5 NDCG@5 Acc@5 NDCG@5 Acc@5 NDCG@5 Acc@5 NDCG@5
STRelay+STGN 0.3604 0.2690 0.4625 0.3412 0.2914 0.2081 0.3841 0.2819
w/o spatial 0.3384 0.2574 0.4467 0.3304 0.2832 0.2039 0.3667 0.2675
w/o temporal 0.3344 0.2547 0.4475 0.3315 0.2793 0.2013 0.3756 0.2743
w/o relaying 0.3418 0.2599 0.4510 0.3332 0.2881 0.2079 0.3697 0.2713
STRelay+Flashback 0.3779 0.2849 0.4803 0.3552 0.3116 0.2243 0.4092 0.3001
w/o spatial 0.3670 0.2782 0.4587 0.3356 0.3075 0.2223 0.3941 0.2894
w/o temporal 0.3686 0.2792 0.4620 0.3381 0.3106 0.2238 0.4044 0.2977
w/o relaying 0.3726 0.2810 0.4728 0.3506 0.3092 0.2235 0.3965 0.2914
STRelay+SNPM 0.3781 0.2846 0.4801 0.3521 0.3279 0.2382 0.4167 0.3017
w/o spatial 0.3682 0.2790 0.4721 0.3464 0.3169 0.2261 0.4101 0.3034
w/o temporal 0.3719 0.2811 0.4710 0.3419 0.3199 0.2279 0.4137 0.3043
w/o relaying 0.3709 0.2793 0.4810 0.3526 0.3162 0.2259 0.4140 0.3039
STRelay+Graph-Flashback 0.3807 0.2873 0.4863 0.3555 0.3348 0.2433 0.4235 0.3113
w/o spatial 0.3737 0.2827 0.4784 0.3526 0.3202 0.2301 0.4180 0.3086
w/o temporal 0.3766 0.2845 0.4764 0.3477 0.3201 0.2292 0.4241 0.3105
w/o 0.3684 0.2787 0.4770 0.3503 0.3331 0.2399 0.4117 0.3025
Experiment Settings

### v2

@@ 4.2.1 Base Location Prediction Model for History Encoding
4.2.2 Multi-Task Location Prediction
After the overall context embedding is generated, it is then fed into a multi-layer perceptron (MLP) followed by a softmax function to predict the next location:
@@ 4.3 Model Training
5 Experiments
TABLE I: Datasets Statistics
@@ 5 Experiments
TABLE II: Performance Comparison of Baselines and STRelay-Enhanced Methods on Istanbul and Tokyo Datasets
Istanbul Tokyo
Methods Acc@5 Acc@10 NDCG@5 NDCG@10 MRR Acc@5 Acc@10 NDCG@5 NDCG@10 MRR
BPR 0.2532 0.3816 0.1821 0.2500 0.1528 0.2133 0.3215 0.1565 0.2121 0.1317
Caser 0.3179 0.3751 0.2464 0.2649 0.2396 0.4115 0.5000 0.2988 0.3275 0.2830
STGCN 0.3312 0.3949 0.2514 0.2721 0.2428 0.4331 0.5306 0.3104 0.3422 0.2927
DeepMove 0.2633 0.3368 0.1903 0.2141 0.1859 0.4127 0.5107 0.3020 0.3338 0.2884
STAN 0.3014 0.3716 0.2240 0.2467 0.2171 0.3911 0.4890 0.2830 0.3147 0.2702
GETNext 0.3107 0.3816 0.2306 0.2536 0.2230 0.4214 0.5197 0.3089 0.3408 0.2948
STGN 0.3292 0.3936 0.2495 0.2704 0.2410 0.4415 0.5342 0.3228 0.3529 0.3056
Flashback 0.3475 0.4139 0.2632 0.2847 0.2557 0.4533 0.5459 0.3339 0.3640 0.3119
Graph-Flashback 0.3521 0.4173 0.2711 0.2925 0.2578 0.4589 0.5548 0.3346 0.3647 0.3154
SNPM 0.3458 0.4125 0.2633 0.2849 0.2527 0.4538 0.5470 0.3305 0.3607 0.3109
LoTNext 0.3530 0.4199 0.2684 0.2901 0.2587 0.4581 0.5534 0.3350 0.3660 0.3165
STRealy+STGN 0.3604 0.4316 0.2690 0.2921 0.2576 0.4625 0.5569 0.3412 0.3709 0.3232
STRealy+Flashback 0.3779 0.4473 0.2847 0.3072 0.2721 0.4803 0.5743 0.3552 0.3859 0.3377
STRealy+SNPM 0.3781 0.4484 0.2846 0.3070 0.2723 0.4801 0.5786 0.3521 0.3837 0.3329
STRealy+LoTNext 0.3744 0.4455 0.2821 0.3051 0.2704 0.4856 0.5828 0.3612 0.3929 0.3423
STRealy+Graph-Flashback 0.3807 0.4525 0.2873 0.3101 0.2743 0.4863 0.5835 0.3555 0.3862 0.3381
TABLE III: Performance Comparison of Baselines and STRelay-Enhanced Methods on Singapore and Moscow Datasets
Singapore Moscow
Methods Acc@5 Acc@10 NDCG@5 NDCG@10 MRR Acc@5 Acc@10 NDCG@5 NDCG@10 MRR
BPR 0.2077 0.3193 0.1466 0.2069 0.1176 0.1362 0.2093 0.0937 0.1273 0.0820
Caser 0.2392 0.3091 0.1726 0.1953 0.1600 0.3239 0.3974 0.2341 0.2579 0.2301
STGCN 0.2917 0.3691 0.2102 0.2353 0.2138 0.3552 0.4237 0.2587 0.2810 0.2437
DeepMove 0.2923 0.3871 0.2053 0.2360 0.2009 0.2905 0.3790 0.2035 0.2322 0.1967
STAN 0.2939 0.3855 0.2086 0.2382 0.2040 0.3105 0.3993 0.2186 0.2474 0.2102
GETNext 0.3171 0.4008 0.2293 0.2564 0.2217 0.3528 0.4302 0.2580 0.2831 0.2456
STGN 0.2736 0.3538 0.1950 0.2209 0.1905 0.3475 0.4157 0.2533 0.2755 0.2392
Flashback 0.2995 0.3815 0.2159 0.2425 0.2092 0.3780 0.4410 0.2802 0.3007 0.2633
Graph-Flashback 0.3207 0.4041 0.2314 0.2584 0.2233 0.3986 0.4645 0.2967 0.3197 0.2765
SNPM 0.3138 0.3993 0.2268 0.2520 0.2164 0.3913 0.4562 0.2906 0.3118 0.2738
LoTNext 0.3241 0.4093 0.2342 0.2618 0.2260 0.4025 0.4719 0.2990 0.3216 0.2814
STRealy+STGN 0.2914 0.3748 0.2080 0.2350 0.2024 0.3841 0.4633 0.2819 0.3076 0.2669
STRealy+Flashback 0.3116 0.3978 0.2243 0.2523 0.2175 0.4009 0.4740 0.2939 0.3177 0.2756
STRealy+SNPM 0.3279 0.4172 0.2382 0.2663 0.2259 0.4167 0.4944 0.3047 0.3288 0.2877
STRealy+LoTNext 0.3312 0.4188 0.2402 0.2686 0.2323 0.4192 0.4949 0.3092 0.3338 0.2905
STRealy+Graph-Flashback 0.3348 0.4251 0.2433 0.2728 0.2338 0.4217 0.4970 0.3113 0.3360 0.2915
5.1 Experiment Setup
5.1.1 Dataset
@@
We evaluate our STRelay using four widely used real-world datasets: Istanbul, Tokyo, Singapore and Moscow. They are collected from April 2012 to February 2013 [33]. Each check-in consists of the user ID, POI ID, latitude, longitude, and timestamp. We follow the data pre-processing steps in [15]. The initial 80% of the data (in chronological order) is used for training, and the remaining data for testing. Table I summarizes the details of the four processed datasets.
5.1.2 Baselines
We choose the following models for location prediction as baselines:
  • BPR [60] employs a pairwise ranking loss to learn user preferences; it exploits direct user-item interaction to separate negative items from positive items to alleviate the sparsity of data and improves the performance.
  • Caser [61] embeds POIs in user interaction history as images, using convolutional filters to capture both local and global sequential dependencies.
  • STGN [10]: add additional gates controlled by the spatiotemporal distances between successive check-ins to LSTM. Note that we add user preference modeling in our experiments for this method.
  • STGCN [10] is a variant of STGN with coupled input and forget gates for improved efficiency. Note that we add user preference modeling in our experiments for this method.
  • DeepMove [57] adds an attention mechanism to GRU for location prediction over sparse mobility traces.
  • STAN [38] models relative spatial-temporal information among POIs with a bi-layer attention architecture.
  • GETNext [21] utilizes a global trajectory flow graph to enhance the prediction of the next POI, alongside proposing a GCN model for generating effective POI embeddings.
  • Flashback [15] is an RNN-based framework that explicitly leverages high-order spatiotemporal distance information to identify informative past hidden states.
  • Graph-Flashback [16] constructs a spatial-temporal knowledge graph for POI representation learning.
  • SNPM [46] represents POI relationships through knowledge graph embeddings and aggregates contextual information from neighboring nodes.
  • LoTNext [48] employs a Long-Tailed Graph Adjustment module to mitigate the impact of long-tailed nodes within the User-POI Interaction Graph.
@@ 5.1.2 Baselines
Among these baselines, we instantiate our STRelay with five techniques: STGN, Flashback, Graph-Flashback, SNPM, and LoTNext to validate its effectiveness.
5.1.3 Evaluation Metrics
@@
We adopt two evaluation metrics commonly used in prior research: Mean Reciprocal Rank (MRR), average Accuracy@K (Acc@K), and Normalized Discounted Cumulative Gain@K (NDCG@K) to evaluate the performance. Their definitions are as follows:
@@
TABLE IV: Performance Comparison of Different STRelay Variants on Istanbul and Tokyo Datasets
Istanbul Tokyo
Methods Acc@5 Acc@10 NDCG@5 NDCG@10 MRR Acc@5 Acc@10 NDCG@5 NDCG@10 MRR
STRelay+STGN 0.3604 0.4316 0.2690 0.2921 0.2576 0.4625 0.5569 0.3412 0.3709 0.3232
w/o spatial 0.3384 0.4050 0.2574 0.2790 0.2491 0.4467 0.5373 0.3304 0.3599 0.3138
w/o temporal 0.3344 0.4013 0.2547 0.2763 0.2468 0.4475 0.5381 0.3315 0.3609 0.3148
w/o relaying 0.3418 0.4095 0.2599 0.2819 0.2515 0.4510 0.5437 0.3332 0.3633 0.3161
w/o spatiotemporal 0.3292 0.3936 0.2495 0.2704 0.2410 0.4415 0.5342 0.3228 0.3529 0.3056
STRealy+Flashback 0.3779 0.4473 0.2849 0.3073 0.2721 0.4803 0.5743 0.3552 0.3859 0.3377
w/o spatial 0.3670 0.4340 0.2782 0.2999 0.2668 0.4587 0.5559 0.3356 0.3672 0.3171
w/o temporal 0.3686 0.4366 0.2792 0.3012 0.2677 0.4620 0.5586 0.3381 0.3695 0.3194
w/o relaying 0.3726 0.4404 0.2810 0.3029 0.2688 0.4728 0.5651 0.3506 0.3807 0.3316
w/o spatiotemporal 0.3475 0.4139 0.2632 0.2847 0.2557 0.4533 0.5459 0.3339 0.3640 0.3119
STRealy+SNPM 0.3781 0.4484 0.2846 0.3070 0.2723 0.4801 0.5786 0.3521 0.3837 0.3329
w/o spatial 0.3682 0.4364 0.2790 0.3011 0.2677 0.4721 0.5695 0.3464 0.3781 0.3271
w/o temporal 0.3719 0.4402 0.2811 0.3032 0.2691 0.4710 0.5715 0.3419 0.3747 0.3222
w/o relaying 0.3709 0.4403 0.2793 0.3018 0.2674 0.4791 0.5776 0.3516 0.3831 0.3327
w/o spatiotemporal 0.3458 0.4125 0.2633 0.2849 0.2527 0.4538 0.5470 0.3305 0.3607 0.3109
STRealy+LoTNext 0.3744 0.4455 0.2821 0.3051 0.2704 0.4856 0.5828 0.3612 0.3929 0.3423
w/o spatial 0.3652 0.4337 0.2772 0.2993 0.2665 0.4733 0.5690 0.3509 0.3821 0.3325
w/o temporal 0.3685 0.4372 0.2789 0.3011 0.2675 0.4746 0.5700 0.3489 0.3800 0.3294
w/o relaying 0.3742 0.4451 0.2817 0.3045 0.2695 0.4829 0.5786 0.3573 0.3885 0.3377
w/o spatiotemporal 0.3530 0.4199 0.2684 0.2901 0.2587 0.4581 0.5534 0.3350 0.3660 0.3165
STRealy+Graph-Flashback 0.3807 0.4525 0.2873 0.3101 0.2743 0.4863 0.5835 0.3555 0.3862 0.3381
w/o spatial 0.3737 0.4432 0.2827 0.3052 0.2709 0.4784 0.5763 0.3526 0.3844 0.3333
w/o temporal 0.3766 0.4453 0.2845 0.3068 0.2723 0.4764 0.5773 0.3477 0.3805 0.3279
w/o relaying 0.3684 0.4368 0.2787 0.3009 0.2672 0.4770 0.5741 0.3503 0.3819 0.3305
w/o spatiotemporal 0.3521 0.4173 0.2711 0.2925 0.2578 0.4589 0.5548 0.3346 0.3647 0.3154
TABLE V: Performance Comparison of Different STRelay Variants on Singapore and Moscow Datasets
Singapore Moscow
Methods Acc@5 Acc@10 NDCG@5 NDCG@10 MRR Acc@5 Acc@10 NDCG@5 NDCG@10 MRR
STRelay+STGN 0.2914 0.3748 0.2081 0.2347 0.2024 0.3841 0.4633 0.2819 0.3076 0.2669
w/o spatial 0.2832 0.3597 0.2039 0.2286 0.1980 0.3667 0.4361 0.2675 0.2901 0.2517
w/o temporal 0.2793 0.3550 0.2013 0.2258 0.1958 0.3756 0.4443 0.2743 0.2966 0.2574
w/o relaying 0.2881 0.3659 0.2079 0.2330 0.2017 0.3697 0.4420 0.2713 0.2948 0.2562
w/o spatiotemporal 0.2736 0.3538 0.1950 0.2209 0.1905 0.3475 0.4157 0.2533 0.2755 0.2392
STRealy+Flashback 0.3116 0.3978 0.2243 0.2523 0.2175 0.4092 0.4861 0.3001 0.3251 0.2816
w/o spatial 0.3075 0.3902 0.2223 0.2491 0.2155 0.3941 0.4688 0.2894 0.3137 0.2723
w/o temporal 0.3106 0.3939 0.2238 0.2508 0.2167 0.4044 0.4764 0.2977 0.3212 0.2794
w/o relaying 0.3092 0.3942 0.2235 0.2511 0.2173 0.3965 0.4707 0.2914 0.3156 0.2740
w/o spatiotemporal 0.2995 0.3815 0.2159 0.2425 0.2092 0.3780 0.4410 0.2802 0.3007 0.2633
STRealy+SNPM 0.3279 0.4172 0.2382 0.2663 0.2259 0.4167 0.4944 0.3047 0.3288 0.2877
w/o spatial 0.3169 0.4040 0.2261 0.2543 0.2185 0.4101 0.4827 0.3034 0.3270 0.2851
w/o temporal 0.3199 0.4115 0.2279 0.2575 0.2206 0.4137 0.4877 0.3043 0.3284 0.2853
w/o relaying 0.3162 0.4015 0.2259 0.2536 0.2182 0.4140 0.4899 0.3039 0.3286 0.2859
w/o spatiotemporal 0.3138 0.3993 0.2268 0.2520 0.2164 0.3913 0.4562 0.2906 0.3118 0.2738
STRealy+LoTNext 0.3312 0.4188 0.2402 0.2686 0.2323 0.4192 0.4949 0.3092 0.3338 0.2905
w/o spatial 0.3274 0.4170 0.2379 0.2670 0.2307 0.4163 0.4897 0.3081 0.3320 0.2894
w/o temporal 0.3264 0.4141 0.2378 0.2662 0.2305 0.4183 0.4948 0.3080 0.3323 0.2881
w/o relaying 0.3277 0.4163 0.2389 0.2676 0.2318 0.4182 0.4927 0.3090 0.3337 0.2903
w/o spatiotemporal 0.3241 0.4093 0.2342 0.2618 0.2260 0.4025 0.4719 0.2990 0.3216 0.2814
STRealy+Graph-Flashback 0.3348 0.4251 0.2433 0.2728 0.2338 0.4235 0.4993 0.3113 0.3360 0.2915
w/o spatial 0.3220 0.4060 0.2322 0.2594 0.2240 0.4180 0.4934 0.3086 0.3331 0.2898
w/o temporal 0.3214 0.4052 0.2319 0.2590 0.2237 0.4241 0.4986 0.3105 0.3349 0.2903
w/o relaying 0.3313 0.4219 0.2392 0.2686 0.2324 0.4117 0.4827 0.3025 0.3256 0.2832
w/o spatiotemporal 0.3207 0.4041 0.2314 0.2584 0.2233 0.3986 0.4645 0.2967 0.3197 0.2765
5.1.4 Experiment Settings